# FlowSymm: Physics Aware, Symmetry Preserving Graph Attention for Network Flow Completion

### v1

@@ 1.2 Our Approach
We introduce FlowSymm, a solver that augments a graph attention backbone with symmetry‑preserving, physics‑aware corrections tailored to partially observed flow graphs. Our formulation is guided by four complementary design principles. (i) First, we reinterpret admissible divergence‑free adjustments as elements of an Abelian group action; by constructing an orthonormal basis and retaining only the leading $k$ basis vectors, we obtain a tractable latent space that scales linearly with the number of missing flows. Unlike typical equivariant GNNs, which embed spatial or permutation symmetries, our group‑action basis directly spans the algebraic space of divergence-free adjustments tailored to the missing flows. (ii) Second, we combine this basis with edge‑wise GATv2 embeddings (Brody et al., 2022): the attention mechanism scores each basis vector in a context‑aware manner, enabling the model to inject corrective flows where and when the local structure demands, while preserving flow balance on the anchor manifold. This attention‑guided selection departs from prior bilevel diagonal regularizers by coupling missing edges through context‑dependent weights over physically interpretable basis actions, rather than treating each edge independently. (iii) Third, we impose soft physical consistency through a feature‑conditioned Tikhonov refinement and train all weights end‑to‑end via reverse‑mode implicit differentiation. (iv) Finally, extensive experiments on three real‑world benchmarks, traffic, power, and bike, show that FlowSymm consistently improves upon nine baselines, reducing RMSE by up to ten percent and yielding attention maps and regularization weights that align with domain‑specific intuition. Taken together, these contributions demonstrate that coupling learned graph attention with symmetry‑preserving flow corrections offers a simple and interpretable route to state‑of‑the‑art network‑flow completion.
@@ 2 Flow Estimation Problem
We lift $\delta^{\mathrm{miss}}\in\mathbb{R}^{m_{\mathrm{miss}}}$ into $\mathbb{R}^{m}$ via $S_{\mathrm{miss}}^{\!\top}\delta^{\mathrm{miss}}$. A noisy observation $\hat{f}\in\mathbb{R}^{m}$ satisfies $\hat{f}_{e}=f_{e}+\varepsilon_{e}$ for $e\in\mathcal{E}_{\mathrm{obs}}$ and $\hat{f}_{e}=0$ on $\mathcal{E}_{\mathrm{miss}}$. We make no explicit assumption on the statistical distribution of the noise; our reconstruction is agnostic to any particular noise model. We abbreviate the measured part as $\hat{f}^{\mathrm{obs}}\coloneqq S_{\mathrm{obs}}\hat{f}$.
Our goal is to reconstruct the missing values $\{f_{e}:e\in\mathcal{E}_{\mathrm{miss}}\}$. We learn a mapping $(\mathcal{G},\hat{f},c)\mapsto\tilde{f}_{\theta}\in\mathbb{R}^{m}$ that (i) fits the noisy measurements (small $\|S_{\mathrm{obs}}\tilde{f}_{\theta}-\hat{f}^{\mathrm{obs}}\|_{2}$), (ii) keeps the balance residual small (small $\|B\tilde{f}_{\theta}-c\|_{2}$), (iii) operates within (and respects) the admissible divergence-free group-action subspace (i.e., the null-space of valid adjustments) that leaves observed flows unchanged, and (iv) couples missing edges through features $X$ rather than a diagonal penalty.
@@
The first estimate respects conservation (i.e., it satisfies $Bf=c$, reducing to divergence-free only when $c=\mathbf{0}$). The final prediction is then allowed a small, learnable violation of node balance so it can handle sensor noise and better fit the measurements, incurring only a small residual relative to $c$. We provide a list of symbols in Appendix A.
@@ 3.2 Group–action basis
where $U\in\mathbb{R}^{m\times k}$ is a column‑orthonormal basis for $\mathcal{A}$. The identity corresponds to $\alpha=0$; composition satisfies $g_{\alpha}\!\circ g_{\beta}=g_{\alpha+\beta}$; and each element is invertible via $g_{\alpha}^{-1}=g_{-\alpha}$. Because every column of $U$ lies in $\ker B$ and vanishes on $\mathcal{E}_{\mathrm{obs}}$, the action preserves both the prescribed injection and the observed flows. To construct $U$ we note that
@@
whose dimension is $r=m_{\mathrm{miss}}-\operatorname{rank}(B_{\mathrm{miss}})\geq m_{\mathrm{miss}}-(n-1)$ for a connected graph (with equality when $B_{\mathrm{miss}}$ has row rank n1n{-}1). In practice we truncate this space and keep only the first $k$ basis vectors to control capacity and compute time; in our experiments we found that retaining $k=256$ basis vectors was sufficient to capture these variations (see Sec. 4.3). Formally, let
@@ 3.3 Attention-guided group-action selection
The term $q_{e,i}$ measures how much the local context of edge $e$ votes for the utilization of basis vector $i$. Each basis vector represents a global adjustment that spans multiple edges. Intuitively, the weight on this basis vector is the sum of the attention scores on its constituent edges. Summing these scores across all missing flows yields a global vector $s\in\mathbb{R}^{k}$, ensuring that the final global correction $\Delta$ is determined by the aggregate demands of local features. Applying a softmax turns $s$ into a probability distribution $\alpha_{\theta}$ over the $k$ basis vectors. Figure 2 shows the histogram of these learned attention weights $\{\alpha_{\theta,i}\}_{i=1}^{k}$ for the Traffic, Power, and Bike datasets (see Sec. 4). Although every basis vector receives some weight, most of the mass concentrates in a few of them, indicating that the model focuses on a subset of group actions.
@@ 3.5 Implicit bilevel training
Inner solve. On each fold, we solve the Tikhonov problem in equation 1 with fold-indexed quantities, producing $\delta^{\mathrm{tik},(k)}_{\theta}$. The associated normal equations mirror §3.4 and are omitted for brevity.
@@ 4 Experiments
We evaluate FlowSymm on three real‐world flow graphs that span transportation, energy and micromobility domains. For every dataset we (i) keep the original directed/undirected topology, (ii) normalize flow magnitudes to $[0,1]$, (iii) encode the features in a one-hot manner, and (iv) follow the 10-fold edge hold-out protocol: in each fold a disjoint subset ${\mathcal{E}}_{\mathrm{obs}}$ is revealed for training, while the complementary ${\mathcal{E}}_{\mathrm{val}}$ is used for validation/testing. For a more detailed description of the datasets and results, please refer to the Appendix C.
@@ 5.1 Limitations
Despite these benefits, there are some limitations. We train on single–shot graphs and do not model temporal dynamics. Extending the approach with recurrent encoders or time–coupled group actions is a promising direction. While $k=256$ sufficed in our graphs, very large or highly cyclic networks may require more basis vectors, increasing memory; adaptive selection strategies could mitigate this. Additionally, we follow a randomized edge hold-out protocol, and if real sensor placements exhibit systematic biases (e.g., clustered in urban centers), performance could degrade under different masking patterns. Furthermore, in cases where unobserved cycles are fully determined by known injections, our basis construction may overstate the degrees of freedom; however, the attention mechanism mitigates this by learning to suppress redundant basis vectors. Lastly, the inner solver is convex, but the overall bilevel training is non-convex and may converge to suboptimal local minima, making hyperparameter initialization and tuning important.
@@ 5.2 Future Work
We use the continental European transmission system as an undirected graph whose vertices are electrical buses and whose edges include both physical transmission lines and auxiliary bus connectors for generation and consumption. Active power flows (in MW) are taken from a single snapshot generated by PyPSA-Eur Brown et al. (2018); Hörsch et al. (2018), a Python toolbox that solves optimal linear power-flow problems using historical load and generation data under default settings. The resulting network comprises 2,048 buses and 2,729 edges. Each edge is equipped with a 14-dimensional feature vector: for line edges we include reactance $x$, resistance $r$, nominal capacity $s_{\mathrm{nom}}$, a binary flag indicating whether $s_{\mathrm{nom}}$ is extendable, the capital cost of capacity expansion, line length, number of parallel circuits, and optimized capacity $s_{\mathrm{nom}}^{\mathrm{opt}}$; for bus edges we encode the control strategy (PQ, PV, or slack) via one-hot encoding. A binary indicator distinguishes line versus bus edges, and all categorical attributes are represented with one-hot vectors Silva et al. (2021). PyPSA-Eur is released under multiple licenses: original source code under the MIT License, documentation under CC-BY-4.0, configuration files mostly under CC0-1.0, and data files under CC-BY-4.0.

### v2

@@ 1.2 Our Approach
We introduce FlowSymm, a solver that augments a graph attention backbone with symmetry‑preserving, physics‑aware corrections tailored to partially observed flow graphs. Our formulation is guided by four complementary design principles. (i) First, we reinterpret admissible divergence‑free adjustments as elements of an Abelian group action; by constructing an orthonormal basis and retaining only the leading $k$ basis vectors, we obtain a tractable latent space that scales linearly with the number of missing flows. Unlike typical equivariant GNNs, which embed spatial or permutation symmetries, our group‑action basis directly spans the algebraic space of divergence-free adjustments tailored to the missing flows. (ii) Second, we combine this basis with edge‑wise GATv2 embeddings (Brody et al., 2022): the attention mechanism scores each basis vector in a context‑aware manner, enabling the model to inject corrective flows where and when the local structure demands, while preserving flow balance on the anchor manifold. This attention‑guided selection departs from prior bilevel diagonal regularizers by coupling missing edges through context‑dependent weights over physically interpretable basis actions, rather than treating each edge independently. (iii) Third, we impose soft physical consistency through a feature‑conditioned Tikhonov refinement and train all weights end‑to‑end via reverse‑mode implicit differentiation. (iv) Finally, we demonstrate through extensive experiments on three real-world benchmarks —traffic, power, and bike— that FlowSymm consistently outperforms nine baselines. Our approach reduces RMSE by up to ten percent while yielding attention maps and regularization weights that align with domain-specific intuition. Taken together, these contributions demonstrate that coupling learned graph attention with symmetry-preserving flow corrections offers a simple and interpretable route to state-of-the-art network-flow completion.
@@ 2 Flow Estimation Problem
We lift $\delta^{\mathrm{miss}}\in\mathbb{R}^{m_{\mathrm{miss}}}$ into $\mathbb{R}^{m}$ via $S_{\mathrm{miss}}^{\!\top}\delta^{\mathrm{miss}}$. A noisy observation $\hat{f}\in\mathbb{R}^{m}$ satisfies $\hat{f}_{e}=f_{e}+\varepsilon_{e}$ for $e\in\mathcal{E}_{\mathrm{obs}}$ and $\hat{f}_{e}=0$ on $\mathcal{E}_{\mathrm{miss}}$. We make no explicit assumption on the statistical distribution of the noise; our reconstruction is agnostic to any particular noise model, and abbreviate the measured part as $\hat{f}^{\mathrm{obs}}\coloneqq S_{\mathrm{obs}}\hat{f}$.
Our goal is to reconstruct the missing values $\{f_{e}:e\in\mathcal{E}_{\mathrm{miss}}\}$. So, we learn a mapping $(\mathcal{G},\hat{f},c)\mapsto\tilde{f}_{\theta}\in\mathbb{R}^{m}$ that (i) fits the noisy measurements (small $\|S_{\mathrm{obs}}\tilde{f}_{\theta}-\hat{f}^{\mathrm{obs}}\|_{2}$), (ii) keeps the balance residual small (small $\|B\tilde{f}_{\theta}-c\|_{2}$), (iii) operates within (and respects) the admissible divergence-free group-action subspace (i.e., the null-space of valid adjustments) that leaves observed flows unchanged, and (iv) couples missing edges through features $X$ rather than a diagonal penalty.
@@
The initial estimate respects conservation (i.e., it satisfies $Bf=c$, reducing to divergence-free only when $c=\mathbf{0}$). The final prediction is then allowed a small, learnable violation of node balance so it can handle sensor noise and better fit the measurements, incurring only a small residual relative to $c$. We provide a list of symbols in Appendix A.
@@ 3.2 Group–action basis
where $U\in\mathbb{R}^{m\times k}$ is a column‑orthonormal basis for $\mathcal{A}$. The identity corresponds to $\alpha=0$; composition satisfies $g_{\alpha}\!\circ g_{\beta}=g_{\alpha+\beta}$; and each element is invertible via $g_{\alpha}^{-1}=g_{-\alpha}$. Because every column of $U$ lies in $\ker B$ and vanishes on $\mathcal{E}_{\mathrm{obs}}$, the action preserves both the prescribed injection and the observed flows. In order to construct $U$, we note that
@@
whose dimension is $r=m_{\mathrm{miss}}-\operatorname{rank}(B_{\mathrm{miss}})\geq m_{\mathrm{miss}}-(n-1)$ for a connected graph (with equality when $B_{\mathrm{miss}}$ has row rank n1n{-}1). In practice we truncate this space and keep only the first $k$ basis vectors to control capacity and compute time; in our experiments, we found that retaining $k=256$ basis vectors was sufficient to capture these variations (see Sec. 4.3). Formally, let
@@ 3.3 Attention-guided group-action selection
The term $q_{e,i}$ measures how much the local context of edge $e$ votes for the utilization of basis vector $i$. Each basis vector represents a global adjustment that spans multiple edges. Intuitively, the weight on this basis vector is the sum of the attention scores on its constituent edges. Summing these scores across all missing flows yields a global vector $s\in\mathbb{R}^{k}$, ensuring that the final global correction $\Delta$ is determined by the aggregate demands of local features. Applying a softmax turns $s$ into a probability distribution $\alpha_{\theta}$ over the $k$ basis vectors. Figure 2 shows the histogram of these learned attention weights $\{\alpha_{\theta,i}\}_{i=1}^{k}$ for the Traffic, Power, and Bike datasets (see Sec. 4). Although every basis vector receives some weight, most of the mass concentrates in a few of them, indicating that the model focuses on a small subset of group actions.
@@ 3.5 Implicit bilevel training
Inner solver. On each fold, we solve the Tikhonov problem in equation 1 with fold-indexed quantities, producing $\delta^{\mathrm{tik},(k)}_{\theta}$. The associated normal equations mirror §3.4 and are omitted for brevity.
@@ 4 Experiments
We evaluate FlowSymm on three real‐world flow graphs that span transportation, energy and micromobility domains. For every dataset we (i) keep the original directed/undirected topology, (ii) normalize flow magnitudes to $[0,1]$, (iii) encode the features in a one-hot manner, and (iv) follow a 10-fold edge hold-out protocol: in each fold a disjoint subset ${\mathcal{E}}_{\mathrm{obs}}$ is revealed for training, while the complementary ${\mathcal{E}}_{\mathrm{val}}$ is used for validation/testing. For a more detailed description of the datasets and results, please refer to the Appendix C.
@@ 5.1 Limitations
Despite the discussed benefits, there are some limitations. We train on single–shot graphs and do not model temporal dynamics. Extending the approach with recurrent encoders or time–coupled group actions is a promising direction. While $k=256$ sufficed in our graphs, very large or highly cyclic networks may require more basis vectors, increasing memory; adaptive selection strategies could mitigate this. Additionally, we follow a randomized edge hold-out protocol, and if real sensor placements exhibit systematic biases (e.g., clustered in urban centers), performance could degrade under different masking patterns. Furthermore, in cases where unobserved cycles are fully determined by known injections, our basis construction may overstate the degrees of freedom; however, the attention mechanism mitigates this by learning to suppress redundant basis vectors. Lastly, the inner solver is convex, but the overall bilevel training is non-convex and may converge to suboptimal local minima, making hyperparameter initialization and tuning important.
@@ 5.2 Future Work
5.3 Acknowledgements
This material is based upon work supported by the National Science Foundation under grant no. 2229876 and is supported in part by funds provided by the National Science Foundation, by the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or its federal agency and industry partners.
@@ 5.3 Acknowledgements
We use the continental European transmission system as an undirected graph whose vertices are electrical buses and whose edges include both physical transmission lines and auxiliary bus connectors for generation and consumption. Active power flows (in MW) are taken from a single snapshot generated by PyPSA-Eur Brown et al. (2018); Hörsch et al. (2018), a Python toolbox that solves optimal linear power-flow problems using historical load and generation data under default settings. The resulting network comprises 2,048 buses and 2,729 edges. Each edge is equipped with a 14-dimensional feature vector: for line edges, we include reactance $x$, resistance $r$, nominal capacity $s_{\mathrm{nom}}$, a binary flag indicating whether $s_{\mathrm{nom}}$ is extendable, the capital cost of capacity expansion, line length, number of parallel circuits, and optimized capacity $s_{\mathrm{nom}}^{\mathrm{opt}}$; for bus edges, we encode the control strategy (PQ, PV, or slack) via one-hot encoding. A binary indicator distinguishes line versus bus edges, and all categorical attributes are represented with one-hot vectors Silva et al. (2021). PyPSA-Eur is released under multiple licenses: original source code under the MIT License, documentation under CC-BY-4.0, configuration files mostly under CC0-1.0, and data files under CC-BY-4.0.